OpenStudy (anonymous):

Rewrite the expression in a form that does not contain absolute values. |x-2| + |x-3|, if 2 < x < 3

hartnn (hartnn):

since x>2, |x-2| >0 so |x-2|=(x-2) since x<3, |x-3| <0 so |x-3|=-(x-3) so what is (x-2)-(x-3) = ?

OpenStudy (anonymous):

if x is between 2 and 3 the \[\left| x-2 \right|\] will open with a positive sign and the \[\left| x-3 \right|\] will open with a negative sign hence \[\left\{(+1)\left( x-2 \right) + (-1)(x-3) \right\}\] hence the final answer will be \[\left\{ x-2 -x+3 \right\} \]

OpenStudy (anonymous):

you can simplify further ;)

OpenStudy (anonymous):

isnt the answer1

hartnn (hartnn):

Expression is x-2-x+3 its simplification = 1

OpenStudy (anonymous):

@hartnn doesnt the expression also become 1, it's constant with no terms involving x, x-2-x+3 would mean nothing as a expression

OpenStudy (anonymous):

I think the analysis done at the top are all good. Bravo to all . . .